Ok, I seriously need help. It's a must, someone please answer this!
There is a line with a slope of 7/3.
What are the points of a line pependicular through the point(8,-1)?
What are the points of a line parrel to the point (8,-1)?
----> if this helps, the slope intercept form is y=7/3x-49/3 & the x and y-intercept coordinates are (0,-49/3).



Answer by scott8148(6628)

the question is a little garbled but seems straightforward

a perpendicular line has a slope that is a negative reciprocal ___ in this case m = -3/7
___ y = (-3/7)x + b
___ substitute the (8,-1) point to find b
___ (-1) = (-3/7)(8) + b ___ -1 + (24/7) = b ___ 17/7 = b
___ so the perpendicular equation is ___ y = (-3/7)x + (17/7) ___ or ___ 7y = -3x + 17

a parallel line has the same slope ___ in this case m = 7/3
___ the solution technique is the same as the perpendicular case